Find the solution for the inequality:
|x/2+4|>5

Found 2 solutions by 4419875, ikleyn:
Answer by 4419875(21)

Find the solution for the inequality:
|x/2+4|>5
Note: |f(x)|>a => f(x) < -a or f(x) > a
two solutions
x/2 + 4 > -5 = x + 8 > -10
Add both sides with 8 to isolate x
= x + 8 - 8 < -10 - 8
= x < -18
x/2 + 4 < 5 = x + 8 < 10
Add both sides with 8 to isolate x
= x + 8 - 8 10 - 8
= x > 2
The solution are x < -18 or x > 2
interval notation -> (-∞, 18) U (2, ∞)

Answer by ikleyn(52777)

.
The solution in the interval notation is (-infinity,-18) U (2,infinity).